# Equivariant and Invariant Parametrized Topological Complexity

### v3

@@ 4. Invariant Parametrized Topological Complexity
$F:=\Psi\circ H\circ(s\times\mathrm{id}_{I})\colon U\times I\to E\times_{B/G}E$
is a $(G\times G)$-homotopy such that $F_{0}=\Psi\circ\mathrm{id}_{E^{I}_{B}\times_{E/G}E^{I}_{B}}\circ s=i_{U}$ and $F_{1}(u)=\Psi(H_{1}(s(u)))=((s_{1}(u))(1),(s_{2}(u))(0))$. As $s(u)=(s_{1}(u),s_{2}(u))\in E^{I}_{B}\times_{E/G}E^{I}_{B}$ for all $u\in U$, it follows $F_{1}(u)=((s_{1}(u))(1),(s_{2}(u))(0))\in E\times_{E/G}E$. Hence, $F_{1}$ is the desired $(G\times G)$-homotopy.
$(2)\implies(1)$. Suppose $H\colon U\times I\to E\times_{B/G}E$ is a $(G\times G)$-homotopy between $f$ and $i_{U}$. Let $s\colon U\to E^{I}_{B}\times_{E/G}E^{I}_{B}$ be the $(G\times G)$-map given by $s(u)=(c_{\pi_{1}(f(u))},c_{\pi_{2}(f(u))})$, where $\pi_{i}\colon E\times_{B/G}E\to E$ is the projection map onto the $i$-th factor. The map $s$ is well-defined since $f$ takes values in $E\times_{E/G}E$. By $G$-homotopy lifting property of $\Psi$, there exists a $(G\times G)$-homotopy $\widetilde{H}\colon U\times I\to E^{I}_{B}\times_{E/G}E^{I}_{B}$ such that the following diagram
[…]
@@ 4.1. Properties and Bounds
[…]
@@
Suppose $Z$ is a topological space with $(G\times G)$-maps $k=(k_{1},k_{2})\colon Z\to E^{I}_{B}\times_{E/G}E^{I}_{B}$ and $h=(h_{1},h_{2})\colon Z\to E^{\prime}\times_{B^{\prime}/G}E^{\prime}$ such that $\Psi\circ k=h$. As $\Psi\circ k=h$, we have
@@
follows from [10, Proposition 4.3]. ∎
@@
Let $p\colon E\to B$ be a $G$-fibration. If $e\in E^{G}$, then the fibre $F=p^{-1}(p(e))$ is a $G$-space and
@@
    if $E\times_{B/G}E$ is a connected $(G\times G)$-$\mathrm{CW}$-complex, then
@@
If $e\in E^{G}$, then $b=p(e)\in B^{G}$. Hence, by Proposition 4.9, $F:=p^{-1}(b)$ admits a $G$-action and $\mathrm{TC}^{G}(F)\leq\mathrm{TC}^{G}[p\colon E\to B].$ Observe that $e\in F^{G}$. Therefore, the inequality $\mathrm{cat}_{G}(F)\leq\mathrm{TC}^{G}(F)$ follows from [5, Proposition 2.7].
(1) Note that if $c_{e}$ is the constant path in $E$ which takes the value $e$, then $(c_{e},c_{e})\in(E^{I}_{B}\times_{E/G}E^{I}_{B})^{(G\times G)}$. Moreover, since $E\times_{B/G}E$ is $(G\times G)$-connected, it follows that
@@
by [27, Corollary 1.12].
Now the last inequality follows from $(1)$ and $(2)$. ∎
@@
Let $p\colon E\to B$ be a $G$-fibration such that $\mathrm{TC}^{G}[p\colon E\to B]=1$. If $e\in E^{G}$, then the fibre $F=p^{-1}(p(e))$ is a $G$-contractible space.
@@
By Proposition 4.10, we have $\mathrm{cat}_{G}(F)=1$, i.e., $F$ is $G$-contractible. ∎
We now establish sufficient conditions for $\mathrm{TC}^{G}[p\colon E\to B]$ to be $1$. This serves as a converse of Corollary 4.11.
@@
Suppose $p\colon E\to B$ is a $G$-fibration such that $E\times_{B/G}E$ is a $G$-$\mathrm{CW}$-complex. Let $e\in E^{G}$. If the fibre $F=p^{-1}(p(e))$ satisfies either
@@
  • $F$ strong $G$-deformation retracts to the point $e$,
@@
=$\mathcal{F}=\{\gamma\in F^{I}\mid\gamma(1/2)=e,\gamma(0)=g\cdot\gamma(1)\},$
@@
Suppose $F$ is $G$-connected, $G$-contractible and $F^{G}=\{e\}$. Since $F$ is $G$-connected, we have ${}_{\{e\}}\mathrm{cat}_{G}(F)=\mathrm{cat}_{G}(F)$, see [25, Remark 2.3] and [10, Lemma 3.14]. Hence, ${}_{\{e\}}\mathrm{cat}_{G}(F)=1$ as $F$ is $G$-contractible. Thus, there exists a $G$-homotopy $H\colon F\times I\to F$ such that $H(f,0)=f$ and $H(f,1)=e$ for all $f\in F$. Let $K\colon F^{I}\times I\to F^{I}$ be the homotopy given by $K(\delta,t)(s)=H(\delta(s),t)$ for all $s,t\in I$ and $\delta\in F^{I}$. Note that $K$ is a $G$-homotopy. If $\gamma\in\mathcal{F}$, then
@@
Suppose $F$ strong $G$-deformation retracts to the point $e$, then there exists a $G$-homotopy $H\colon F\times I\to F$ such that $H(f,0)=f$ and $H(f,1)=e$ and $H(e,t)=e$ for all $f\in F$ and $t\in I$. Then the homotopy $K$ defined on $F^{I}$ like above satisfies $K(\gamma,t)(1/2)=e$ due to the condition $H(e,t)=e$ for all $t\in I$.
@@
Later, we will also provide sufficient conditions for $\mathrm{TC}^{G}[p\colon E\to B]=1$ and its converse when the group action on the base is free, as stated in Corollary 4.31 and Corollary 4.15, respectively.
@@
Suppose $\Psi_{K}\colon E^{I}_{B}\times_{E/K}E^{I}_{B}\to E\times_{B/K}E$ is the invariant parametrized fibration corresponding to $K$-fibration $p$. Then the following diagram
[…]
@@
where $F$ is the fibre of $p$.
@@
Note that, by [20, Theorem 3], the map $p\colon E\to B$ is a $K$-fibration. Hence, the result follows from [8, Page 235] and Proposition 4.13. ∎
Corollary 4.15.
Suppose $G$ is a compact Hausdorff topological group and $p\colon E\to B$ is a $G$-fibration with fibre $F$ such that $G$ acts freely on $B$. If $\mathrm{TC}^{G}[p\colon E\to B]=1$, then the $F$ is contractible.
Proof.
This follows from Corollary 4.14 and the fact that $\mathrm{TC}(F)=1$ if and only if $F$ is contractible. ∎
@@ 4.1.1. Cohomological Lower Bounds
where $\gamma^{\prime}_{t}(s)=\gamma(s+t(1-s))$ and $\delta^{\prime}_{t}(s)=\delta(s(1-t))$. Then following the proof of Theorem 4.7, we see that $H$ is well-defined, $(G\times G)$-equivariant, $H(\gamma,\delta,0)=(\gamma,\delta)$, and $H(\gamma,\delta,1)=(c_{\gamma(1)},c_{\delta(0)})$. Hence, $c\circ f$ is $(G\times G)$-homotopic to the identity map of $E^{I}_{B}\times_{E/G}E^{I}_{B}$. ∎
@@
[…]
@@
Note that $\Psi\circ c=i$ implies $(c^{h}_{G^{2}})^{*}\circ(\Psi^{h}_{G^{2}})^{*}=(i^{h}_{G^{2}})^{*}$. Since $c$ is a $(G\times G)$-homotopy equivalence (see Lemma 4.16), it follows $c^{h}_{G^{2}}$ is a homotopy equivalence. Hence, $(c^{h}_{G^{2}})^{*}$ is an isomorphism. Thus, the result follows from Theorem 2.2. ∎
Remark 4.18.
We note that any $G$-map $f\colon X\to Y$ that is a non-equivariant homotopy equivalence induces an isomorphism on the level of Borel cohomology, see [28]. Hence, for Theorem 4.17, we don’t need $c$ is a $(G\times G)$-homotopy equivalence, we only require $c$ to be a $(G\times G)$-map and a non-equivariant homotopy equivalence.
Now we give a non-equivariant cohomological lower bound for the invariant parametrized topological complexity. Let $E_{B,G^{2}}:=(E\times_{B/G}E)/(G\times G)$ and let $\mathbb{k}_{G^{2}}E\subseteq E_{B,G^{2}}$ denote the image of the saturated diagonal subspace $\mathbb{k}(E)=E\times_{E/G}E\subseteq E\times_{B/G}E$ under the orbit map $\rho\colon E\times_{B/G}E\to E_{B,G^{2}}$. By using Theorem 4.7 and following the arguments in Theorem 3.6, one can establish the following theorem. The proof is left to the reader.
@@ 4.1.2. Product Inequalities
by [1, Proposition 3.7]. ∎
The product inequality for invariant topological complexity was proved in [25, Theorem 3.18]. In the following corollary, we show that the cofibration hypothesis assumed in [25, Theorem 3.18] can be removed by using Theorem 4.20.
@@
by Proposition 4.4 and Theorem 4.20. ∎
The proof of the following corollary is similar to Corollary 3.8 and can be shown using Corollary 4.14 and Theorem 4.20.
@@
Suppose $p_{i}\colon E_{i}\to B_{i}$ is a $G$-fibration such that $G$ acts on $B_{i}$ freely for $i=1,2$. If $G$ is compact Hausdorff, then $p_{1}\times p_{2}\colon E_{1}\times E_{2}\to B_{1}\times B_{2}$ is a $G$-fibration, where $G$ acts diagonally on the spaces $E_{1}\times E_{2}$ and $B_{1}\times B_{2}$. Furthermore, if $E_{1}$ and $E_{2}$ are Hausdorff, and $E_{1}\times E_{1}\times E_{2}\times E_{2}$ is completely normal, then
@@
The proof of the following corollary is similar to that of Corollary 3.9 and follows from Proposition 4.9 and Corollary 4.22.
@@
Suppose $p_{i}\colon E_{i}\to B$ is a $G$-fibration, for $i=1,2$, such that $G$ acts on $B$ freely. Let $E_{1}\times_{B}E_{2}=\{(e_{1},e_{2})\in E_{1}\times E_{2}\mid p_{1}(e_{1})=p_{2}(e_{2})\}$ and let $p\colon E_{1}\times_{B}E_{2}\to B$ be the $G$-map given by $p(e_{1},e_{2})=p_{1}(e_{1})=p_{2}(e_{2})$, where $G$ acts on $E_{1}\times_{B}E_{2}$ diagonally. If $G$ is compact Hausdorff, then $p$ is a $G$-fibration. Furthermore, if $E_{1}$ and $E_{2}$ are Hausdorff, and $E_{1}\times E_{1}\times E_{2}\times E_{2}$ is completely normal, then
@@ 4.2. Some technical results
In this subsection, we establish two technical results which will help us compute the invariant parametrized topological complexity of Fadell-Neuwirth fibrations in Section 5.
Definition 4.24 ([20, Section 5]).
@@
As noted in Theorem 2.5, the induced map $\overline{p}\colon\overline{E}\to\overline{B}$ is a $G$-fibration when $G$ is a compact Hausdorff topological group. However, to compute the invariant parametrized topological complexity of the equivariant Fadell-Neuwirth fibration, defined in [12], we require Proposition 4.25, which says $\overline{p}$ is also locally trivial. We will introduce equivariant Fadell-Neuwirth fibration and calculate their invariant parametrized topological complexity in Section 5. Now, we present one more result which will be required in Section 5.

### v4

@@ 4. Invariant Parametrized Topological Complexity
$F:=\Psi\circ H\circ(\sigma\times\mathrm{id}_{I})\colon U\times I\to E\times_{B/G}E$
is a $(G\times G)$-homotopy such that $F_{0}=\Psi\circ\mathrm{id}_{E^{I}_{B}\times_{E/G}E^{I}_{B}}\circ\sigma=i_{U}$ and $F_{1}(u)=\Psi(H_{1}(\sigma(u)))=((\sigma_{1}(u))(1),(\sigma_{2}(u))(0))$. As $\sigma(u)=(\sigma_{1}(u),\sigma_{2}(u))\in E^{I}_{B}\times_{E/G}E^{I}_{B}$ for all $u\in U$, it follows $F_{1}(u)=((\sigma_{1}(u))(1),(\sigma_{2}(u))(0))\in E\times_{E/G}E$. Hence, $F_{1}$ is the desired $(G\times G)$-homotopy.
$(2)\implies(1)$. Suppose $H\colon U\times I\to E\times_{B/G}E$ is a $(G\times G)$-homotopy between $f$ and $i_{U}$. Let $\sigma\colon U\to E^{I}_{B}\times_{E/G}E^{I}_{B}$ be the $(G\times G)$-map given by $\sigma(u)=(c_{\pi_{1}(f(u))},c_{\pi_{2}(f(u))})$, where $\pi_{i}\colon E\times_{B/G}E\to E$ is the projection map onto the $i$-th factor. The map $\sigma$ is well-defined, since $f$ takes values in $E\times_{E/G}E$. By the $G$-homotopy lifting property of $\Psi$, there exists a $(G\times G)$-homotopy $\widetilde{H}\colon U\times I\to E^{I}_{B}\times_{E/G}E^{I}_{B}$ such that the following diagram
$U\times\{0\}$$E^{I}_{B}\times_{E/G}E^{I}_{B}$$U\times I$$E\times_{B/G}E$$\scriptstyle{\sigma}$$\scriptstyle{\Psi}$$\scriptstyle{H}$$\scriptstyle{\widetilde{H}}$
@@ 4.1. Properties and bounds
$(E^{\prime})^{I}_{B^{\prime}}\times_{E^{\prime}/G}(E^{\prime})^{I}_{B^{\prime}}$$E^{I}_{B}\times_{E/G}E^{I}_{B}$$E^{\prime}\times_{B^{\prime}/G}E^{\prime}$$E\times_{B/G}E,$$\scriptstyle{\Psi^{\prime}}$$\scriptstyle{\Psi}$
@@
Suppose $Z$ is a topological space with $(G\times G)$-maps $k=(k_{1},k_{2})\colon Z\to E^{I}_{B}\times_{E/G}E^{I}_{B}$ and $h=(h_{1},h_{2})\colon Z\to E^{\prime}\times_{B^{\prime}/G}E^{\prime}$ such that $\Psi\circ k=h$. Then we have
@@
follows from [colmangranteqtc, Proposition 4.3]. ∎
@@
Suppose $G$ is a compact Hausdorff topological group, and let $p\colon E\to B$ be a $G$-fibration. If $e\in E^{G}$, then the fibre $F=p^{-1}(p(e))$ is a $G$-space, and
@@
    if $G$ is a compact Lie group and $E\times_{B/G}E$ is a connected $(G\times G)$-$\mathrm{CW}$-complex, then
@@
If $e\in E^{G}$, then $b=p(e)\in B^{G}$. Hence, by Proposition˜4.8, $F:=p^{-1}(b)$ admits a $G$-action and $\mathrm{TC}^{G}(F)\leq\mathrm{TC}^{G}[p\colon E\to B].$ Observe that $e\in F^{G}$. Therefore, the inequality $\mathrm{cat}_{G}(F)\leq\mathrm{TC}^{G}(F)$ follows from [ZK, Proposition 2.7].
(1) Note that if $c_{e}$ is the constant path in $E$ which takes the value $e\in E$, then $(c_{e},c_{e})\in(E^{I}_{B}\times_{E/G}E^{I}_{B})^{(G\times G)}$. Moreover, since $E\times_{B/G}E$ is $(G\times G)$-connected, it follows that
@@
by [M, Corollary 1.12]. Now the last inequality follows from $(1)$ and $(2)$. ∎
@@
Suppose $G$ is a compact Hausdorff topological group, and let $p\colon E\to B$ be a $G$-fibration such that $\mathrm{TC}^{G}[p\colon E\to B]=1$. If $e\in E^{G}$, then the fibre $F=p^{-1}(p(e))$ is a $G$-contractible space.
@@
By Proposition˜4.9, we have $\mathrm{cat}_{G}(F)=1$, i.e., $F$ is $G$-contractible. ∎
We now establish sufficient conditions for $\mathrm{TC}^{G}[p\colon E\to B]$ to be $1$. This serves as a converse of Corollary˜4.10.
@@
Suppose $G$ is a compact Hausdorff topological group. Suppose $p\colon E\to B$ is a $G$-fibration such that $E\times_{B/G}E$ is a $G$-$\mathrm{CW}$-complex. Let $e\in E^{G}$. If the fibre $F=p^{-1}(p(e))$ satisfies either
@@
  • $F$ strongly $G$-deformation retracts to the point $e$,
@@
$\mathcal{F}=\{\gamma\in F^{I}\mid\gamma(1/2)=e,\gamma(0)=g\cdot\gamma(1)\},$
@@
Suppose $F$ is $G$-connected, $G$-contractible and $F^{G}=\{e\}$. Since $F$ is $G$-connected, we have $_{\{e\}}\mathrm{cat}_{G}(F)=\mathrm{cat}_{G}(F)$, see [invarianttc, Remark 2.3] and [colmangranteqtc, Lemma 3.14]. Hence, $_{\{e\}}\mathrm{cat}_{G}(F)=1$ as $F$ is $G$-contractible. Thus, there exists a $G$-homotopy $H\colon F\times I\to F$ such that $H(f,0)=f$ and $H(f,1)=e$ for all $f\in F$. Let $K\colon F^{I}\times I\to F^{I}$ be the homotopy given by $K(\delta,t)(s)=H(\delta(s),t)$ for all $s,t\in I$ and $\delta\in F^{I}$. Note that $K$ is a $G$-homotopy. If $\gamma\in\mathcal{F}$, then
@@
Suppose $F$ strongly $G$-deformation retracts to the point $e$, then there exists a $G$-homotopy $H\colon F\times I\to F$ such that $H(f,0)=f$ and $H(f,1)=e$ and $H(e,t)=e$ for all $f\in F$ and $t\in I$. Then the homotopy $K$ defined on $F^{I}$ as above satisfies $K(\gamma,t)(1/2)=e$ due to the condition $H(e,t)=e$ for all $t\in I$.
@@
Later, we will also provide sufficient conditions for $\mathrm{TC}^{G}[p\colon E\to B]=1$ and its converse when the group action on the base is free, as stated in Corollary˜4.28 and Corollary˜4.13, respectively.
@@
Suppose $\Psi_{K}\colon E^{I}_{B}\times_{E/K}E^{I}_{B}\to E\times_{B/K}E$ is the invariant parametrized fibration corresponding to $K$-fibration pp. Then the following diagram
$E^{I}_{B}\times_{E/K}E^{I}_{B}$$E^{I}_{B}\times_{E/G}E^{I}_{B}$$E\times_{B/K}E$$E\times_{B/G}E$$\scriptstyle{\Psi_{K}}$$\scriptstyle{\Psi}$
@@
where $F$ denotes the fibre of $p$. Moreover, if $\mathrm{TC}^{G}[p\colon E\to B]=1$, then $F$ is contractible.
@@
Note that, by [gevorgyan2023equivariant, Theorem 3], the map $p\colon E\to B$ is a $K$-fibration. Hence, the desired inequalities follow from Proposition˜4.12 and [farber-para-tc, Page 235]. The last statement follows from $\mathrm{TC}(F)=1$ if and only if $F$ is contractible, see [FarberTC, Theorem 1]. ∎
@@ 4.1.1. Cohomological Lower Bounds
where $\gamma^{\prime}_{t}(s)=\gamma(s+t(1-s))$ and $\delta^{\prime}_{t}(s)=\delta(s(1-t))$. Then following the proof of Theorem˜4.6, we see that $H$ is well-defined, $(G\times G)$-equivariant, $H(\gamma,\delta,0)=(\gamma,\delta)$, and $H(\gamma,\delta,1)=(c_{\gamma(1)},c_{\delta(0)})$. Hence, $c\circ f$ is $(G\times G)$-homotopic to the identity map of $E^{I}_{B}\times_{E/G}E^{I}_{B}$. ∎
@@
$E\times_{E/G}E$$E^{I}_{B}\times_{E/G}E^{I}_{B}$$E\times_{B/G}E$$\scriptstyle{c}$$\scriptstyle{i}$$\scriptstyle{\Psi}$
@@
Note that $\Psi\circ c=i$ implies $(c^{h}_{G^{2}})^{*}\circ(\Psi^{h}_{G^{2}})^{*}=(i^{h}_{G^{2}})^{*}$. Since $c$ is a $(G\times G)$-homotopy equivalence (see Lemma˜4.14), it follows $c^{h}_{G^{2}}$ is a homotopy equivalence. Hence, $(c^{h}_{G^{2}})^{*}$ is an isomorphism. Thus, the result follows from Theorem˜2.7. ∎
Now we give a non-equivariant cohomological lower bound for the invariant parametrized topological complexity. Let $E_{B,G^{2}}:=(E\times_{B/G}E)/(G\times G)$ and let $\mathbb{k}_{G^{2}}E\subseteq E_{B,G^{2}}$ denote the image of the saturated diagonal subspace $\mathbb{k}(E)=E\times_{E/G}E\subseteq E\times_{B/G}E$ under the orbit map $\rho\colon E\times_{B/G}E\to E_{B,G^{2}}$. By using Theorem˜4.6 and following the arguments in Theorem˜3.6, one can establish the following theorem. The proof is left to the reader.
@@ 4.1.2. Product Inequalities
by [A-D-S, Proposition 3.7]. ∎
The product inequality for invariant topological complexity was proved in [invarianttc, Theorem 3.18]. In the following corollary, we show that the cofibration hypothesis assumed in [invarianttc, Theorem 3.18] can be removed by using Theorem˜4.17.
@@
by Proposition˜4.4 and Theorem˜4.17. ∎
The proof of the following corollary is similar to Corollary˜3.9 and can be shown using Corollary˜4.13 and Theorem˜4.17.
@@
Suppose $p_{i}\colon E_{i}\to B_{i}$ is a $G$-fibration such that $G$ acts on $B_{i}$ freely for $i=1,2$. If $G$ is compact Hausdorff, then $p_{1}\times p_{2}\colon E_{1}\times E_{2}\to B_{1}\times B_{2}$ is a $G$-fibration, where $G$ acts diagonally on the spaces $E_{1}\times E_{2}$ and $B_{1}\times B_{2}$. Furthermore, if $E_{1}\times E_{1}\times E_{2}\times E_{2}$ is completely normal, then
@@
The proof of the following corollary is similar to that of Corollary˜3.10 and follows from Proposition˜4.8 and Corollary˜4.19.
@@
Suppose $p_{i}\colon E_{i}\to B$ is a $G$-fibration, for $i=1,2$, such that $G$ acts on $B$ freely. Let $E_{1}\times_{B}E_{2}=\{(e_{1},e_{2})\in E_{1}\times E_{2}\mid p_{1}(e_{1})=p_{2}(e_{2})\}$ and let $p\colon E_{1}\times_{B}E_{2}\to B$ be the $G$-map given by $p(e_{1},e_{2})=p_{1}(e_{1})=p_{2}(e_{2})$, where $G$ acts on $E_{1}\times_{B}E_{2}$ diagonally. If $G$ is compact Hausdorff, then $p$ is a $G$-fibration. Furthermore, if $E_{1}\times E_{1}\times E_{2}\times E_{2}$ is completely normal, then
@@ 4.2. Some technical results
In this subsection, we establish two technical results which will help us compute the invariant parametrized topological complexity of Fadell-Neuwirth fibrations in Section˜5.
Definition 4.21.
@@
As noted in Theorem˜2.10, the induced map $\overline{p}\colon\overline{E}\to\overline{B}$ is a $G$-fibration when $G$ is a compact Hausdorff topological group. However, to compute the invariant parametrized topological complexity of the equivariant Fadell-Neuwirth fibration, defined in [D-EqPTC], we require Proposition˜4.22, which says $\overline{p}$ is also locally trivial. We will introduce equivariant Fadell-Neuwirth fibrations and calculate their invariant parametrized topological complexity in Section˜5. Now, we present one more result which will be required in Section˜5.